OpenStudy (anonymous):

The sum of four consecutive integers is 74. What is the first integer? 16 17 18 19

OpenStudy (phi):

call the unknown first number x four consecutive integers means integers "next" to each other, like 1,2,3,4 or 9,10,11,12 notice that each one is 1 bigger than the one before it. If we start with the x as the first number, the next one up will be x+1 the one after that will be (x+1) + 1 (or x+2) and the last one x+3 all 4 add up to 74: x + (x+1) + (x+2) + (x+3) = 74 can you solve this for x ?

OpenStudy (the_fizicx99):

4x + 6 = 74 -6 -6 4x = 68 /4 /4 x = 17. 17 + 18 + 19 + 20 = 74 So the first would be 17